OpenStudy (anonymous):

which property is used in the equation mg + mh = m(g+h) associative commutative distibutive identity?

OpenStudy (anonymous):

distributive. aka. a(b+c) = ab+ac. associative is (a+b) + c = a + (b+c). commutative is a+b+c = c+b+a. order doesnt matter additive identity = a+0=a.

OpenStudy (anonymous):

Well it looks like associative but backwords?